# Neutrino Heating in 1D, 2D, and 3D core-collapse supernovae: characterizing the explosion of high-compactness stars

### v1

@@ 4 Interplay between neutrino heating and convection
Figure 4: Integrated turbulent energy in the gain region (left panel) and neutrino heating in the gain region (middle panel) for 1D+ and 1D simulations of the progenitors listed in Table 1. The integrated turbulent energy is roughly half of the neutrino heating. The dashed lines in the two rightmost panels show the semi-analytical expression in Eq. (14), which is only valid during the stalled-shock phase, and therefore we only compute it starting at 100 ms, until either the shock crosses 500 km or the simulation ends. The excellent agreement between dashed and solid lines shows that Eq. (14) is a good description of neutrino heating in both 1D and 1D+ simulations, i.e. with or without convection.
To carry out this part of the analysis, we compare the same nine 1D and 1D+ simulations from Boccioli (2024a) already discussed in Section 1. The agreement between the semi-analytical expression in Eq. (14) and the neutrino heating calculated directly from the simulation is excellent, for both 1D and 1D+ simulations, as shown in Figure 4 by the solid and dashed lines. It should be highlighted that the difference between the two rightmost panels of Figure 4 is only the inclusion of $\nu$-driven convection, with everything else in the simulation being kept the same. For completeness, we show in the left panel the rate of turbulent energy generation in the gain region $E_{\rm g}$, defined in Eq. (9), which is roughly half of the neutrino heating, although the exact value depends on compactness. If one considers neutrino heating plus turbulent energy generation, this is compatible with the findings of (Gogilashvili et al., 2023a) that convection decreases the explosion condition by $\sim$ 30 %.
@@
Figure 7: Neutrino root mean squared energies (upper panel) and neutrino luminosities calculated at 500 km (lower panel) for electron-type neutrinos. These are 1D (dashed lines) and 1D+ (solid lines) simulations for the progenitors listed in Table 1. The luminosities are very similar since they mostly depend on the accretion history, which are the same in 1D and 1D+ (but also, to very good approximation, multi-D) simulations. Deviations only appear after the explosion sets in, when the mass accretion rate and temperature of the neutrinosphere significantly drop. A similar argument holds for neutrino energies. Neutrino energies for electron antineutrinos have a very similar time dependence and are $\sim 10\%$ larger. Luminosities, on the other hand, are quite similar in the accretion phase, i.e. later than $\sim 100$ ms after bounce.
@@
Figure 8: The vertical axis shows the time at which $\dot{Q}_{\nu}^{\rm max}$ is reached. For the left panel, the horizontal axis shows the explosion time, which is defined as the time when the shock crosses 500 km. The right panels have on the horizontal axis the time when the Si/O interface is accreted through the shock. The middle one is color-coded according to compactness, whereas the right panel is color-coded according to explosion outcome (i.e. successful or failed). Solid black lines in all panels show the bisectors, i.e. $t_{\rm max}=t_{\rm Si/O}$. Exploding progenitors show a correlation between $t_{\rm Si/O}$ and $t_{\rm max}$, . Outliers and failed SNe are discussed in the text. There is instead no apparent correlation between $t_{\rm expl}$ and $t_{\rm max}$. This is because high-compactness progenitors have very large explosion energies, and therefore very high shock velocities, as discussed in the text.
@@ 4.1 Interplay among relevant timescales of the problem
It is instructive to analyze the interplay between relevant timescales of the problem, namely the time at which the maximum of $\dot{Q}_{\nu}$ is reached ($t_{\rm max}$), the time when the explosion sets in ($t_{\rm expl}$, defined as the time when shock reaches 500 km), and finally the time when the Si/O interface is accreted ($t_{\rm Si/O}$). The latter is important since it plays an important role in triggering the explosion (Vartanyan et al., 2018; Summa et al., 2016; Lentz et al., 2015) and can be used to predict the outcome of the explosion (Boccioli et al., 2023; Wang et al., 2022; Tsang et al., 2022; Ertl et al., 2016). For this part of the analysis, we will adopt the same 341 1D+ models used by Boccioli (2024a).
@@
It is also instructive to analyze the correlation between $t_{\rm expl}$ and $t_{\rm Si/O}$ shown in Figure 9. For $\xi_{2.0}\lesssim 0.5$, $t_{\rm expl}$ positively correlates with $t_{\rm Si/O}$. For $\xi_{2.0}\gtrsim 0.5$, however, there is no correlation anymore, indicating that the accretion of the Si/O interface is not important for the explosion, in line with the findings of Boccioli (2024b). The density drop at the Si/O interface is important to determine the explodability of progenitors with $\xi_{2.0}<0.5$, but it has no impact on the explodability of progenitors with $\xi_{2.0}>0.5$ that instead all explode, as highlighted by Boccioli & Fragione (2024) and shown again in this work.
@@
Figure 9: Only successful explosions of the 341 1D+ simulations from Boccioli & Fragione (2024) are shown in this Figure. We show the region where $t_{\rm Si/O}>t_{\rm expl}$ in grey, to highlight that there the Si/O interface is accreted after the explosion sets in. Therefore, despite the negative correlation, the accretion of the Si/O interface does not influence the explosion. As discussed in the text, only in the region where $t_{\rm expl}$ positively correlates with $t_{\rm Si/O}$ the accretion of the Si/O interface significantly affects the explosion.
@@
Finally, there are roughly $\sim 80$ progenitors with a wide range of $t_{\rm Si/O}$ and $t_{\rm expl}$, which however all have $t_{\rm max}\approx 100$ ms. As can be seen from the middle panel of Figure 8, these are all low-compactness progenitors, which, as pointed out in Section 3, do not experience the typical increase in $\dot{Q}_{\nu}$ seen in for high-compactness progenitors. Instead, after peaking at $\approx 100$ ms, neutrino heating does not increase significantly during the stalled-shock phase because the temperature of the neutrinosphere does not increase significantly). Therefore, one expects the neutrino heating for these low-compactness progenitors to peak at around 100 ms, which is also confirmed by the 20 3D Fornax simulations analyzed in this paper. Then, some of them will explode after the accretion of a significant density drop at the Si/O interface, and others will not. A more detailed analysis of the interplay between the explosion condition and the accretion of the Si/O interface is given in Boccioli (2024b).
@@ 5 Comparison between STIR and other 1D explosion models
Historically, extensive explodability studies involving hundreds of simulations have been mostly carried out using 1D simulations, with possibly only one exception of a very large 2D study by Nakamura et al. (2015). Notably, in that work, the authors carried out hundreds of 2D simulations and found that most high-compactness progenitors did not explode. The reason for the disagreement in explodability with the 2D simulations presented here is not easy to assess. Possible explanations could be that, in the study of Nakamura et al. (2015), the simulations were run with the IDSA approximate neutrino transport from Liebendörfer et al. (2009) and a simple leakage scheme for heavy-lepton neutrinos. Moreover, they adopted the LS220 nuclear EOS from Lattimer & Swesty (1991), which typically leads to a faster contraction of the PNS compared to the SFHo EOS adopted for most simulations analyzed in this work. The faster PNS contraction typically leads to earlier shock revival (Boccioli et al., 2022; Janka & Bauswein, 2022), but at the same time, it also causes a faster collapse into a black-hole (Schneider et al., 2020). Therefore, this could potentially inhibit the successful explosion of high-compactness progenitors, which tend to explode later, as shown in Figure 8. It should also be noted that, in their simulations, Nakamura et al. (2015) adopt fully Newtonian gravity (i.e. without any general relativistic effective potential), which is known to delay even further (i.e. later than 1 second post bounce) the explosion of high-compactness progenitors (O’Connor & Couch, 2018). Therefore, without any general relativistic corrections, high-compactness progenitors will be harder to explode, which might explain the lack of exploding high-compactness progenitors in the study of Nakamura et al. (2015). As we have shown in previous sections, the time dependence of neutrinosphere temperatures (i.e. neutrino energies) is very important in order to obtain successful shock revival of high-compactness progenitors. Therefore, less accurate neutrino transport schemes such as the IDSA, lack of general relativistic corrections and, more importantly, the simpler leakage scheme for heavy lepton neutrinos, could significantly alter the cooling of the PNS.
@@ 5.1 O’Connor 2011
An often-used explodability criterion, because of its simplicity, is the one developed by O’Connor & Ott (2011). It is based on the compactness $\xi_{2.5}$ calculated at bounce (the difference with the pre-supernova compactness is however rather small), and they found that progenitors with $\xi_{2.5}\lesssim 0.45$ are more likely to explode, whereas the high compactness progenitors are more likely to collapse. Their study has the benefit of exploring a larger set of progenitor models generated with different codes, i.e. both KEPLER (Woosley et al., 2002) and FRANEC (Limongi & Chieffi, 2006) progenitors, as well as different equations of state. However, it was carried out with a simpler version of the code GR1D adopted in this work, with a simple leakage scheme for neutrino transport and where the explosion was achieved by multiplying the neutrino heating term in the leakage scheme by a factor $f_{\rm heat}$ of order 1.
@@
Another well-known problem with artificially enhancing heating in 1D is that unphysical oscillations of the shock radius by hundreds of kilometers occur (Gogilashvili et al., 2023b). Moreover, it is not clear whether the same value of $f_{\rm heat}$ should be applied to all progenitors. As mentioned above, O’Connor & Ott (2011) assumed that multi-D simulations could achieve a fixed heating efficiency independent of compactness, which however is not guaranteed.
@@ 5.2 PUSH
$\mathcal{G}(t)$ is initially zero, and then at a time $t_{\rm on}$ gradually increases until it reaches its maximum value $k_{\rm push}$, and eventually gradually decreases back to zero after the explosion has been launched. As can be seen from Eq. (16), the extra heating in PUSH depends on the luminosity and energy of the heavy lepton neutrinos, which are evolved with the simplified IDSA formalism (Liebendörfer et al., 2009), that is known to not properly capture the time evolution of the neutrino energy. However, the energy dissipation due to $\nu$-driven convection can be directly related to the luminosity of the electron flavor neutrinos (Murphy et al., 2013) or, equivalently, to the neutrino heating in the gain region (Gogilashvili et al., 2023a). This is the first major difference between PUSH and STIR.
@@
Another difference between the PUSH models and the 1D+ (Boccioli et al., 2023; Couch et al., 2020) and 2D models (Vartanyan & Burrows, 2023) discussed in this paper, is the fate of lower mass stars at solar metallicity evolved with the KEPLER stellar evolution code in the mass range $12M_{\odot}\lesssim M_{\rm ZAMS}\lesssim 15M_{\odot}$. PUSH obtains successful explosions for these progenitors whereas (Boccioli et al., 2023; Couch et al., 2020) and (Vartanyan & Burrows, 2023) obtain a failed-SN, with the 1D+ models from Boccioli et al. (2023); Couch et al. (2020) finding even more failed-SN, which might be connected to the discussion in Section 3. This discrepancy could be due to a combination of a few factors.
@@
Moreover, the progenitor ZAMS mass of 1987A is not very well constrained and could be anywhere between 18 and 21 $M_{\odot}$ (or even larger), which for the KEPLER models used in the calibration of PUSH corresponds to values of $\xi_{2.0}$ anywhere between 0.15 and 0.45. Given the uncertainties in the calibration of SN 1987A, one could change the calibration of $k_{\rm push}$ enough to have an overall smaller maximum value of $k_{\rm push}$, located at larger compactnesses which could in principle reproduce the multi-dimensional failed explosions of these KEPLER progenitors with $12M_{\odot}\lesssim M_{\rm ZAMS}\lesssim 15M_{\odot}$. Of course, this could only be verified by a thorough study, where the PUSH calibration is systematically changed (for example with the goal of reproducing Figure 3). The purpose of this study is however only to highlight the differences among the main 1D explosion models. These differences are fundamentally due to different prescriptions of how the extra-heating of the 1D model depends on the pre-supernova properties (i.e. compactness) as well as on the dynamical properties of the explosion (i.e. neutrino luminosity, average energy, etc…).
@@ 5.4 Muller 2016
Another approach to the explodability is the semi-analytical model of Müller et al. (2016a), which does not require complex simulations at all. It is built upon the theoretical framework developed over the years by Janka (2001) and Janka (2012), and it expands it and refines it in several way, also thanks to the significant improvement of self-consistent multi-dimensional simulations. The explodability derived by this approach for KEPLER progenitors is relatively similar to the one derived by Ertl et al. (2016), but also not too dissimilar to the one derived by O’Connor & Ott (2011). Differently from the 2D and 3D results of Vartanyan & Burrows (2023) and the 1D+ results of Boccioli et al. (2023); Couch et al. (2020), progenitors with large compactness tend to yield failed explosions, whereas low-compactness stars are more likely to explode, with a few exceptions for stars around $15M_{\odot}$. According to Müller et al. (2016a), most stars in the range $12M_{\odot}\lesssim M_{\rm ZAMS}\lesssim 15M_{\odot}$ explode, whereas stars in the range $22M_{\odot}\lesssim M_{\rm ZAMS}\lesssim 25M_{\odot}$ and $M_{\rm ZAMS}\gtrsim 30M_{\odot}$ do not.
Interestingly, some of the stars with compactnesses $0.3\lesssim\xi_{2.5}\lesssim 0.4$ (which roughly corresponds to $0.5\lesssim\xi_{2.0}\lesssim 0.7$) experience a successful shock revival that however does not turn into an explosion since the central PNS quickly accretes enough material to form a black hole and therefore the neutrino heating source that is still powering the explosion shuts off, causing the explosion to fail. This has been shown by recent simulations to be indeed a possible scenario. However, despite an early (i.e. $\sim 1.5$ s after bounce) formation of a black hole, one can still achieve relatively moderately to very powerful explosions (Burrows et al., 2023; Chan et al., 2018; Sykes & Müller, 2024; Eggenberger Andersen et al., 2024; Burrows et al., 2024a). Therefore, even though high-compactness progenitors will more quickly form black holes, they can still produce a successful explosion.

### v2

@@ 4 Interplay between neutrino heating and convection
Figure 4: Integrated turbulent energy generation rate in the gain region (left panel) and neutrino heating in the gain region (middle panel) for 1D+ and 1D simulations of the progenitors listed in Table 1. The integrated turbulent energy is roughly half of the neutrino heating. The dashed lines in the two rightmost panels show the semi-analytical expression in Eq. (15), which is only valid during the stalled-shock phase, and therefore we only compute it starting at 100 ms, until either the shock crosses 500 km or the simulation ends. The excellent agreement between dashed and solid lines shows that Eq. (15) is a good description of neutrino heating in both 1D+ and 1D simulations, i.e. with or without convection.
To carry out this part of the analysis, we compare the same nine 1D and 1D+ simulations from Boccioli & Fragione (2024) already discussed in Section 2.2. The agreement between the semi-analytical expression in Eq. (15) and the neutrino heating calculated directly from the simulation is excellent, for both 1D and 1D+ simulations, as shown in Figure 4 by the solid and dashed lines. It should be highlighted that the difference between the two rightmost panels of Figure 4 is only the inclusion of $\nu$-driven convection, with everything else in the simulation being kept the same. For completeness, we show in the left panel the rate of turbulent energy generation in the gain region $E_{\rm g}$, defined in Eq. (9), which is roughly half of the neutrino heating, although the exact value depends on compactness. If one considers neutrino heating plus turbulent energy generation, this is compatible with the findings of (Gogilashvili et al., 2024) that convection decreases the explosion condition by $\sim$ 30 %.
@@
Figure 7: Neutrino root mean squared energies (upper panel) and neutrino luminosities calculated at $R=\langle R\rangle_{g}$ (lower panel) for electron-type neutrinos. These are 1D (dashed lines) and 1D+ (solid lines) simulations for the progenitors listed in Table 1. The luminosities are very similar since they mostly depend on the accretion history, which are the same in 1D and 1D+ (but also, to very good approximation, multi-D) simulations. Deviations only appear after the explosion sets in, when the mass accretion rate and temperature of the neutrinosphere significantly drop. A similar argument holds for neutrino energies. Neutrino energies for electron antineutrinos have a very similar time dependence and are $\sim 10\%$ larger. Luminosities, on the other hand, are quite similar in the accretion phase, i.e. later than $\sim 100$ ms after bounce.
@@
Figure 8: The vertical axis shows the time at which $\dot{Q}_{\nu}^{\rm max}$ is reached. For the left panel, the horizontal axis shows the explosion time, which is defined as the time when the shock crosses 500 km. The right panels have on the horizontal axis the time when the Si/O interface is accreted through the shock. The middle one is color-coded according to compactness, whereas the right panel is color-coded according to explosion outcome (i.e. successful or failed). Solid black lines in all panels show the bisectors, i.e. $t_{\rm max}=t_{\rm Si/O}$, and the arrows in the first panels have sole purpose of guiding the eye. Exploding progenitors show a correlation between $t_{\rm Si/O}$ and $t_{\rm max}$, . Outliers and failed SNe are discussed in the text. There is instead no apparent correlation between $t_{\rm expl}$ and $t_{\rm max}$. This is because high-compactness progenitors have very large explosion energies, and therefore very high shock velocities, as discussed in the text.
@@ 4.1 Interplay among relevant timescales of the problem
It is instructive to analyze the interplay between relevant timescales of the problem, namely the time at which the maximum of $\dot{Q}_{\nu}$ is reached ($t_{\rm max}$), the time when the explosion sets in ($t_{\rm expl}$, defined as the time when shock reaches 500 km), and finally the time when the Si/O interface is accreted ($t_{\rm Si/O}$). The latter is important since it plays an important role in triggering the explosion (Lentz et al., 2015; Summa et al., 2016; Vartanyan et al., 2018) and can be used to predict the outcome of the explosion (Ertl et al., 2016; Wang et al., 2022; Tsang et al., 2022; Boccioli et al., 2023). For this part of the analysis, we will adopt the same 341 1D+ models used by Boccioli & Fragione (2024).
@@
It is also instructive to analyze the correlation between $t_{\rm expl}$ and $t_{\rm Si/O}$ shown in Figure 9. For $\xi_{2.0}\lesssim 0.5$, $t_{\rm expl}$ positively correlates with $t_{\rm Si/O}$. For $\xi_{2.0}\gtrsim 0.5$, however, there is no correlation anymore, indicating that the accretion of the Si/O interface is not important for the explosion, in line with the findings of Boccioli et al. (2025). The density drop at the Si/O interface is important to determine the explodability of progenitors with $\xi_{2.0}<0.5$, but it has no impact on the explodability of progenitors with $\xi_{2.0}>0.5$ that instead all explode, as highlighted by Boccioli & Fragione (2024) and shown again in this work.
@@
Figure 9: Only successful explosions of the 341 1D+ simulations from Boccioli & Fragione (2024) are shown in this Figure. The dashed line indicates where $t_{\rm Si/O}=t_{\rm expl}$. Therefore, points to the right of that refer to simulations for which the Si/O interface is accreted after the explosion sets in. Therefore, despite the negative correlation seen for those points, the accretion of the Si/O interface does not actually influence the explosion. As discussed in the text, only in the region where $t_{\rm expl}$ positively correlates with $t_{\rm Si/O}$ (i.e. left of the dashed line) the accretion of the Si/O interface significantly affects the explosion.
@@
Finally, there are roughly $\sim 80$ progenitors with a wide range of $t_{\rm Si/O}$ and $t_{\rm expl}$, which however all have $t_{\rm max}\approx 100$ ms. As can be seen from the middle panel of Figure 8, these are all low-compactness progenitors, which, as pointed out in Section 3, do not experience the typical increase in $\dot{Q}_{\nu}$ seen in for high-compactness progenitors. Instead, after peaking at $\approx 100$ ms, neutrino heating does not increase significantly during the stalled-shock phase because the temperature of the neutrinosphere does not increase significantly). Therefore, one expects the neutrino heating for these low-compactness progenitors to peak at around 100 ms, which is also confirmed by the 20 3D Fornax simulations analyzed in this paper. Then, some of them will explode after the accretion of a significant density drop at the Si/O interface, and others will not. A more detailed analysis of the interplay between the explosion condition and the accretion of the Si/O interface is given in Boccioli et al. (2025).
@@ 5 Comparison between STIR and other 1D explosion models
Historically, extensive explodability studies involving hundreds of simulations have been mostly carried out using 1D simulations, with possibly only one exception of a very large 2D study by Nakamura et al. (2015). Notably, in that work, the authors carried out hundreds of 2D simulations and found that most high-compactness progenitors did not explode. The reason for the disagreement in explodability with the 2D simulations presented here is not easy to assess. Possible explanations could be that, in the study of Nakamura et al. (2015), the simulations were run with the IDSA approximate neutrino transport from Liebendörfer et al. (2009) and a simple leakage scheme for heavy-lepton neutrinos. Moreover, they adopted the LS220 nuclear EOS from Lattimer & Swesty (1991), which typically leads to a faster contraction of the PNS compared to the SFHo EOS adopted for most simulations analyzed in this work. The faster PNS contraction typically leads to earlier shock revival (Boccioli et al., 2022; Janka & Bauswein, 2023), but at the same time, it also causes a faster collapse into a black-hole (Schneider et al., 2020). Therefore, this could potentially inhibit the successful explosion of high-compactness progenitors, which tend to explode later, as shown in Figure 8. It should also be noted that, in their simulations, Nakamura et al. (2015) adopt fully Newtonian gravity (i.e. without any general relativistic effective potential), which is known to delay even further (i.e. later than 1 second post bounce) the explosion of high-compactness progenitors (O’Connor & Couch, 2018). Therefore, without any general relativistic corrections, high-compactness progenitors will be harder to explode, which might explain the lack of exploding high-compactness progenitors in the study of Nakamura et al. (2015). As we have shown in previous sections, the time dependence of neutrinosphere temperatures (i.e. neutrino energies) is very important in order to obtain successful shock revival of high-compactness progenitors. Therefore, less accurate neutrino transport schemes such as the IDSA, lack of general relativistic corrections and, more importantly, the simpler leakage scheme for heavy lepton neutrinos, could significantly alter the cooling of the PNS.
@@ 5.1 O’Connor 2011
An often-used explodability criterion, because of its simplicity, is the one developed by O’Connor & Ott (2011). It is based on the compactness $\xi_{2.5}$ calculated at bounce (the difference with the pre-supernova compactness is however rather small), and they found that progenitors with $\xi_{2.5}\lesssim 0.45$ are more likely to explode, whereas the high compactness progenitors are more likely to collapse. Their study has the benefit of exploring a larger set of progenitor models generated with different codes, i.e. both KEPLER (Woosley et al., 2002) and FRANEC (Limongi & Chieffi, 2006) progenitors, as well as different equations of state. However, it was carried out with a simpler version of the code GR1D adopted in this work, with a simple leakage scheme to describe neutrino heating, rather than solving for neutrino transport, as it is done in the more recent version of the code adopted in this work through an M1 scheme. Moreover, the explosion was achieved by multiplying the neutrino heating term in the leakage scheme by a factor $f_{\rm heat}$ of order 1.
@@
Another well-known problem with artificially enhancing heating in 1D is that unphysical oscillations of the shock radius by hundreds of kilometers occur (Gogilashvili et al., 2023). Moreover, it is not clear whether the same value of $f_{\rm heat}$ should be applied to all progenitors. As mentioned above, O’Connor & Ott (2011) assumed that multi-D simulations could achieve a fixed heating efficiency independent of compactness, which however is not guaranteed.
@@ 5.2 PUSH
$\mathcal{G}(t)$ is initially zero, and then at a time $t_{\rm on}$ gradually increases until it reaches its maximum value $k_{\rm push}$, and eventually gradually decreases back to zero after the explosion has been launched. As can be seen from Eq. (17), the extra heating in PUSH depends on the luminosity and energy of the heavy lepton neutrinos, which are evolved with the simplified IDSA formalism (Liebendörfer et al., 2009), that is known to not properly capture the time evolution of the neutrino energy. However, the energy dissipation due to $\nu$-driven convection can be directly related to the luminosity of the electron flavor neutrinos (Murphy et al., 2013) or, equivalently, to the neutrino heating in the gain region (Gogilashvili et al., 2024). This is the first major difference between PUSH and STIR.
@@
Another difference between the PUSH models and the 1D+ (Couch et al., 2020; Boccioli et al., 2023) and 2D models (Vartanyan & Burrows, 2023) discussed in this paper, is the fate of lower mass stars at solar metallicity evolved with the KEPLER stellar evolution code in the mass range $12M_{\odot}\lesssim M_{\rm ZAMS}\lesssim 15M_{\odot}$. PUSH obtains successful explosions for these progenitors whereas (Couch et al., 2020; Boccioli et al., 2023) and (Vartanyan & Burrows, 2023) obtain a failed-SN, with the 1D+ models from Boccioli et al. (2023); Couch et al. (2020) finding even more failed-SN, which might be connected to the discussion in Section 3. This discrepancy could be due to a combination of a few factors.
@@
Moreover, the progenitor ZAMS mass of 1987A is not very well constrained and could be anywhere between 18 and 21 $M_{\odot}$ (or even larger), which for the KEPLER models used in the calibration of PUSH corresponds to values of $\xi_{2.0}$ anywhere between 0.15 and 0.45. Given the uncertainties in the calibration of SN 1987A, one could change the calibration of $k_{\rm push}$ enough to have an overall smaller maximum value of $k_{\rm push}$, located at higher compactnesses which could in principle reproduce the multi-dimensional failed explosions of these KEPLER progenitors with $12M_{\odot}\lesssim M_{\rm ZAMS}\lesssim 15M_{\odot}$. Of course, this could only be verified by a thorough study, where the PUSH calibration is systematically changed (for example with the goal of reproducing Figure 3). The purpose of this study is however only to highlight the differences among the main 1D explosion models. These differences are fundamentally due to different prescriptions of how the extra-heating of the 1D model depends on the pre-supernova properties (i.e. compactness) as well as on the dynamical properties of the explosion (i.e. neutrino luminosity, average energy, etc…).
@@ 5.4 Muller 2016
Another approach to the explodability is the semi-analytical model of Müller et al. (2016a), which does not require complex simulations at all. It is built upon the theoretical framework developed over the years by Janka (2001) and Janka (2012), and it expands it and refines it in several way, also thanks to the significant improvement of self-consistent multi-dimensional simulations. The explodability derived by this approach for KEPLER progenitors is relatively similar to the one derived by Ertl et al. (2016), but also not too dissimilar to the one derived by O’Connor & Ott (2011). Differently from the 2D and 3D results of Vartanyan & Burrows (2023) and the 1D+ results of Couch et al. (2020); Boccioli et al. (2023), progenitors with large compactness tend to yield failed explosions, whereas low-compactness stars are more likely to explode, with a few exceptions for stars around $15M_{\odot}$. According to Müller et al. (2016a), most stars in the range $12M_{\odot}\lesssim M_{\rm ZAMS}\lesssim 15M_{\odot}$ explode, whereas stars in the range $22M_{\odot}\lesssim M_{\rm ZAMS}\lesssim 25M_{\odot}$ and $M_{\rm ZAMS}\gtrsim 30M_{\odot}$ do not.
Interestingly, some of the stars with compactnesses $0.3\lesssim\xi_{2.5}\lesssim 0.4$ (which roughly corresponds to $0.5\lesssim\xi_{2.0}\lesssim 0.7$) experience a successful shock revival that however does not turn into an explosion since the central PNS quickly accretes enough material to form a black hole and therefore the neutrino heating source that is still powering the explosion shuts off, causing the explosion to fail. This has been shown by recent simulations to be indeed a possible scenario (Chan et al., 2018; Burrows et al., 2023; Sykes & Müller, 2025; Eggenberger Andersen et al., 2025; Burrows et al., 2024a). However, despite an early (i.e. $\sim 1.5$ s after bounce) formation of a black hole, one can still achieve relatively moderate to very powerful explosions. Therefore, even though high-compactness progenitors will more quickly form black holes, they can still produce a successful explosion.